OpenStudy (anonymous):

a) give an example of a polynomial p(x) without a zero (a zero is a number c such that p(c)=0) b) Give an example of a function h(x) whose graph is above the x-axis and below the x-axis, yet h(x) does not intersect the x-axis by example i'm assuming sketch a graph?

OpenStudy (agent47):

a) x^2+1

OpenStudy (agent47):

2) y = Sec(x)

OpenStudy (agent47):

2) * h(x)=Sec(x)